OpenStudy (anonymous):

what formula is needed to locate the average rate of change for f(x)=x^3 + 2x^2 - 5x - 6. Locate the average rate of change when x = -2 and x = 2.

OpenStudy (anonymous):

Just plug in x as -2 and 2? 2 8+8-10-6 = 0 -2 -8+8+10-6 = 4 Average rate of change is 0 and 4?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Calculus, right?

OpenStudy (anonymous):

algebra ii

OpenStudy (anonymous):

Is f(x) the actual rate of change equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i have 3 minutes left, so please help me out!

OpenStudy (anonymous):

He did.

OpenStudy (anonymous):

the answer is 4?

OpenStudy (anonymous):

He gave you two answers.

OpenStudy (anonymous):

I am pretty sure you were only required to plug in the stated information which was given, if not then I apologize for any error.

OpenStudy (anonymous):

i just really need to get this right.

OpenStudy (anonymous):

and i am completely confused with all of these questions

OpenStudy (anonymous):

Oh, try Aice Math, the confusion with explode beyond explanation. I shall help as much as possible, I am currently doing some easy Physics homework, so won't bother me too much.

OpenStudy (anonymous):

Reminder that the site is used to explain the answer and help you figure out how to do it yourself. If you've got a question, ask.